OpenStudy (anonymous):

help please! A. Line m contains the points, A(–2, 6) and B(4, 8), while line n contains the C(8, 12) and D(x, 24). Given m and n are perpendicular lines, solve for the value of x. Given m and n are parallel lines, solve for the value of x. In your final answer, include all formulas and calculations necessary to solve for x. *question closed but i still need help!*

OpenStudy (anonymous):

@amistre64 can you help ?? im not sure what equations to use

OpenStudy (amistre64):

we are looking to compare slopes, what is the slope formula?

OpenStudy (amistre64):

we also need to know the properties of parallel slopes and perpendicular slopes ...

OpenStudy (anonymous):

m = y2-y1 / x2-x1 is slope right?

OpenStudy (amistre64):

thats the formula, not find the slope of each line

OpenStudy (amistre64):

**now find the slope ....

OpenStudy (anonymous):

for line M i got 1/3 for the slope

OpenStudy (anonymous):

im confused on line N because one of the variables is just X so i dont know what to do

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

Line m contains the points, A(–2, 6) and B(4, 8), (–2, 6) -(4, 8) ------ -6,-2 slope = 1/3 i agree while line n contains the C(8, 12) and D(x, 24). (x, 24) -(8, 12) ---------- x-8, 12 slope is just 12/(x-8) now we have to know the properties of perp and parallel slopes

OpenStudy (amistre64):

parallel slopes are equal, so equate them 1/3 = 12/(x-8) x-8 = 36 what is x? ------------------------------------ perp slopes have a product equal to -1, so set it up 1/3 * 12/(x-8) = -1 12/(x-8) = -3 x-8 = -4 what is x?